FlyLord69winkwink:

differentiate the function y = 1/x. Express the derivative using Leibniz notation

FlyLord69winkwink:

@JSVSL7 please help :)

FlyLord69winkwink:

do i replace the x from F(x+h)-F(x)/h with 1/x?

FlyLord69winkwink:

what do you mean by "how did you call for me"? sorry, i don't get it

FlyLord69winkwink:

so this is what i mean ...one second.

FlyLord69winkwink:

f(x+h)-f(x)/h = f(1/4+h)-f(1/4)all over h

FlyLord69winkwink:

correct me if i'm wrong... the first derivative represents the avg speed, the second derivative represents the acceleration and what does the third derivative rep?

FlyLord69winkwink:

oh i thought the 3rd derivative might mean something ...hmm anyways. thanks

NAGC:

Then y = x √ x − 4 √ x = x 3 2 − 4 x 1 2 Differentiate using the power rule. d y d x = ( 3 2 ) x 1 2 − 2 x − 1 2 = ( 3 2 ) x 1 2 − 2 x 1 2 = ( 3 √ x 2 ) − 2 √ x Get a common denominator of 2 √ x , and you'll arrive at their answer.

NAGC:

there

NAGC:

I just learned this in advanced

NAGC:

did I do it right?

Narad:

The definition of the deivative of a function f(x) according to Leibniz is \[f'(x)=\lim _{h \rightarrow 0} \frac{ f(x+h) -f(x)}{ _{} h}\] Here \[f(x)= 1/x\] You can continue.......